# Eeg-Defender: Defending against Jailbreak through Early Exit Generation of Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly attracting attention in various applications. Nonetheless, there is a growing concern as some users attempt to exploit these models for malicious purposes, including the synthesis of controlled substances and the propagation of disinformation. In an effort to mitigate such risks, the concept of "Alignment" technology has been developed. However, recent studies indicate that this alignment can be undermined using sophisticated prompt engineering or adversarial suffixes, a technique known as "Jailbreak." Our research takes cues from the human-like generate process of LLMs. We identify that while jailbreaking prompts may yield output logits similar to benign prompts, their initial embeddings within the model’s latent space tend to be more analogous to those of malicious prompts. Leveraging this finding, we propose utilizing the early transformer outputs of LLMs as a means to detect malicious inputs, and terminate the generation immediately. Built upon this idea, we introduce a simple yet significant defense approach called Eeg-Defender for LLMs. We conduct comprehensive experiments on ten jailbreak methods across three models. Our results demonstrate that Eeg-Defender is capable of reducing the Attack Success Rate (ASR) by a significant margin, roughly 85% in comparison with 50% for the present SOTAs, with minimal impact on the utility and effectiveness of LLMs.
Warning: this paper may contain offensive prompts and model outputs.
Eeg-Defender: Defending against Jailbreak through Early Exit Generation of Large Language Models
Chongwen Zhao, Zhihao Dou, Kaizhu Huang Duke Kunshan University {chongwen.zhao, zhihao.dou, kaizhu.huang}@dukekunshan.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) are garnering unprecedented attention and application in the field of artificial intelligence, with chatbots such as ChatGPT Achiam et al. (2023) and Llama Touvron et al. (2023a) standing out as notable examples. However, an inherent challenge arises due to the fact that these models could generate inappropriate and potentially harmful content, including biased, unlawful, pornographic, and fraudulent material Weidinger et al. (2021). To mitigate the risks associated with such content and to steer LLM-generated responses away from these issues, researchers have innovated a series of alignment algorithms Ouyang et al. (2022); Wei et al. (2022); Song et al. (2024). Through the implementation of these algorithms, chatbots have been empowered to discern and tactfully refuse to generate outputs in response to prompts that naively seek to elicit potentially harmful content.
(a) Jailbreak embedding visualization (see more details in Section 3.1).
@@
More recently, it has been however discovered that well-designed jailbreak prompts can circumvent such alignment, posing new challenges for building stricter safety barriers Zou et al. (2023); Liu et al. (2024); Wei et al. (2024). Meanwhile, efforts to defend against jailbreaks are ongoing. Prompt-based methods Zhang et al. (2024); Xie et al. (2023); Jain et al. (2023); Wei et al. (2023); Inan et al. (2023a) approach defense by manipulating or detecting user prompts. However, these methods are impractical since they degrade significantly in utility Xu et al. (2024a). As a result, researchers turn to decoding-based defense methods Robey et al. (2024); Cao et al. (2024); Xu et al. (2024a); Zhao et al. (2024b). Instead of directly accessing prompts, decoding-based defense methods leverage the model’s internal properties. Since these methods can maintain high model functionality, decoding-based defense methods have shown promise in defending against jailbreak attacks.
@@
Unfortunately, current decoding-based defense technologies are insufficient. Studies show that present defense methods could only reduce the Attack Success Rate (ASR) by around 50% against jailbreak prompts Xu et al. (2024b). Approaches like RA-LLM Cao et al. (2024) and Smooth-LLM Robey et al. (2024) propose generating responses multiple times with random dropouts to defend against character-sensitive adversarial suffix attacks. However, they are less effective against prompt crafting attacks, which typically involve character-insensitive prompts. SafeDecoding Xu et al. (2024a) aims to increase the likelihood of disclaimer generation artificially, but in practice, it fails to effectively reduce ASR in models with stronger safety barriers.
@@
In response to the drawbacks of existing decoding-based defense methods, we revisit the functions of different layers in LLMs. Todd et al. (2024) reveal that the initial layers specialize in triggering specific tasks. The middle layers act as repositories of knowledge and shape the emotional tone of the output Zhou et al. (2024); Zhao et al. (2024a). Subsequent layers are where the refinement of the language output occurs Fan et al. (2024). Given that language only affects how we deliver, but not the semantics of expression Fedorenko et al. (2024), we postulate that LLMs process jailbreak and harmful prompts similarly when recognizing functions in the initial layers and accessing stored knowledge in the middle layers.
@@
To validate our postulation, we conduct a series of analysis. First, our results in Section 3.2 demonstrate that the classifiers trained on the initial layers achieve over 80% accuracy in detecting fail-to-refuse harmful prompts. More intuitively, as illustrated in Figure 1(a), our empirical visualization shows that starting from the early layers of models (e.g., layer 6 and layer 8), embeddings of jailbreak prompts aligned with harmful prompts. In the middle layers (e.g., layer 12), where LLMs retrieve information, jailbreak embeddings shift towards benign embeddings slightly, and by the later layers (e.g., layers 28 and 32), they become increasingly aligned with benign embeddings. Ultimately, the jailbreak embeddings are either distributed throughout the space (as seen with Llama2) or distributed with the decision boundary (as seen with Vicuna and Guanaco), complicating the model in recognizing jailbreak status.
@@
Remarkably, the process by which large language models generate responses closely mirrors how humans organize language. To structure language output, humans first form an idea Piaget (1926), then draw upon experiences and memories Corballis (2019); Tulving et al. (1972). Finally, language serves as a conduit for conveying information Brandt (2010); Fedorenko et al. (2024). As such, we argue that the focus may be placed on the early or intermediate layers rather than the latter or even final layers, which are overemphasized by current defense methods.
@@
  • Defend jailbreak through early exit. Building on our insights into LLM jailbreak, we propose Eeg-Defender. Eeg-Defender reduces Attack Success Rate (ASR) by approximately 85% against existing jailbreak methods, with near-zero computational cost.
@@ 2.2 LLM Jailbreak
Prompt Crafting. Wei et al. (2024) found that LLMs are often vulnerable to jailbreaks due to competing objectives and mismatched generalizations. They proposed 30 jailbreak methods to elicit harmful responses from GPT and Claude. To reduce the manual effort involved in crafting prompts, Yu et al. (2024); Mehrotra et al. (2024); Chao et al. (2024) developed several automatic frameworks for jailbreaking LLMs. These frameworks typically create a virtual context and suppress the denying output, which utilize the result founded in Wei et al. (2024).
Token Optimizing. In a white-box setting, attackers have access to the gradients of LLMs, allowing them to optimize prompts to increase the likelihood of generating affirmative responses. Zou et al. (2023) achieved jailbreak by optimizing an adversarial suffix to minimize the loss of the desired prefix of outputting. The AutoDAN attack constructed prompts that can pass perplexity testing Liu et al. (2024). Additionally, Qiang et al. (2024) combined In-Context Learning (ICL) with model gradients to distract the model’s attention and generate harmful content.
@@ 2.3 Jailbreak Defense
Prompt-based Defense. Directly detecting content within prompts can help prevent harmful content generated by LLMs. Therefore, Inan et al. (2023a), OpenAI (2023b), and Jigsaw (2017) have proposed several APIs for content detection. In addition to filtering harmful prompts, manipulation of prompts can be incorporated to reinforce safety measures. Zhang et al. (2024) proposed adding prompts that instruct the model to prioritize safety. Xie et al. (2023) leveraged psychological principles by incorporating self-reminder prompts in system messages, encouraging LLMs to respond responsibly and thereby reducing the success rate of jailbreak attacks. Additionally, Jain et al. (2023) outlined three defensive strategies: perplexity detection, paraphrasing, and reorganization. However, this approach suffers from a high false positive rate, limiting its effectiveness in real applications.
Decoding-based Defense. Some jailbreak prompts can be highly sensitive to character-level changes. Therefore, introducing random perturbations and dropouts can help mitigate attack effects Robey et al. (2024). Cao et al. (2024) developed RA-LLM, which leverages the inherent robustness of LLMs and applies Monte Carlo sampling with dropout as a defense strategy. Xu et al. (2024a) revealed that safety disclaimers often remain among the top tokens in the outputs generated by jailbreak prompts. They proposed amplifying these safety token probabilities to reduce the risk of jailbreaks. Besides, Zhao et al. (2024b) identified several safety-critical layers within LLMs and re-aligned these layers to improve overall safety. Overall, these defense methods effectively balance utility and safety, but their effectiveness diminishes with models that have stronger safety barriers.
@@ 2.4 Language Production
One of the most widely accepted theories about how language is organized in humans is Piaget’s theory, which suggests that thought forms first, and then language develops Piaget (1926). When individuals have a concept in mind, they draw upon their memories Corballis (2019) and personal experiences Tulving et al. (1972); Sherwood (2015). Conversely, language is optimized for communication, where people use signs to express and share their thoughts with others; this system of signs has gradually evolved into complex languages Brandt (2010). In summary, language is often seen as a bridge between communication and cognition in humans, with ideas forming first and language being structured based on memories and experiences.
Our work is inspired by the process of language production, a phenomenon also reflected in LLMs. After receiving a prompt, the LLM first identifies the purpose of the prompt and triggers a function within the model Todd et al. (2024). Then, it accesses and processes stored information Meng et al. (2022) and manages emotional tone Zhao et al. (2024a); Zhou et al. (2024) for prompts in the early and middle layers. Several studies found that by truncating Fan et al. (2024), skipping Elhoushi et al. (2024), and pruning Men et al. (2024) some deeper layers, models can respond faster while maintaining correctness. This observation reveals that later layers are responsible for organizing languages. Due to the shared semantic similarities between jailbreak and harmful prompts, we believe that LLMs tend to perform similarly when identifying functions and accessing information.
@@ 3 A Closer Look into Jailbreak
Although concurrent work Lin et al. (2024) demonstrates that well-aligned LLMs can effectively distinguish between benign and harmful prompts within the model’s latent space, the mechanisms behind jailbreaks remain under debate. To gain a deeper understanding of jailbreak, we further investigate the representation of prompts.
Motivated by the human-like generation process of the language model and the observation that well-aligned LLMs can reject malicious and some jailbreak prompts, our aim is to understand how jailbreak prompts manage to bypass safety barriers. Previous attack methods Zou et al. (2023); Wei et al. (2024) suggest that the first token of response influenced the overall responses. Rejection responses always start with an apology or a disclaimer, while helpful responses to benign prompts typically begin with an affirmation. Given that jailbreak prompts share semantic similarities with harmful prompts but resemble benign prompts in their response patterns, we first conjecture that jailbreak embeddings progressively transit from harmful to benign as the layers go deeper.
@@ 3.1 Embedding of Jailbreak: A Toy Example
We conduct a toy example to examine how jailbreak prompts are positioned in the embedding space. We collected 60 benign prompts from Alpaca Eval Li et al. (2023b), and 60 harmful prompts from AdvBench Zou et al. (2023). Then, we evaluated 60 prompts generated by GCG Zou et al. (2023), AutoDAN Liu et al. (2024), GPTFuzz Yu et al. (2024), and Tap Mehrotra et al. (2024), all of which are effective at jailbreak models. As Figure 1(a) shows, in the final layer, the harmful prompts and benign prompts embedding are linearly separable after PCA, with jailbreak embeddings positioned between them, making detection and defense against jailbreaks more challenging. However, we found that in the earlier layers of LLMs (e.g., layer 6), embeddings for benign and harmful prompts are clearly separated, with jailbreak embeddings more closely aligned with harmful prompts. As we move to deeper layers, although benign and harmful embeddings remain distinct, jailbreak embeddings incline toward the center of benign embeddings. With this intriguing phenomenon, we also hypothesize that early and middle layers of transformers inherently possess the ability to discern jailbreak prompts.
@@ 3.2 Shallow Layers can Distinguish Jailbreak
To simulate real-world chatbot applications, we adapted the toxic-chat training dataset Lin et al. (2023) to validate our hypothesis. The dataset includes 5,082 user prompts from the Vicuna online demo, with 384 identified as harmful. Specifically, we re-evaluated the harmful prompts in the dataset using Llama and Vicuna. We identified 302 harmful prompts for Llama2 and 140 harmful prompts for Vicuna that model can successfully reject using keyword matching. Second, we collected the embedding of all layers for benign prompts and rejected harmful prompts and trained 32 MLP classifiers as well as 32 prototype classifiers corresponding to the output of each layer, respectively. We use these two classifier sets to identify jailbreak prompts that the model cannot reject.
As shown in Figure 2, classifiers collected from the early layers perform much better than those from the later layers. The accuracy in distinguishing jailbreak prompts exceed 80% for both models up to the twelfth layer, strongly supporting our second hypothesis. This indicates that we should likely focus on the early and intermediate layer space rather than the output space.
@@
(a) Accuracy on Vicuna
(b) Accuracy on Llama
Figure 2: The accuracy of MLP and prototype classifiers in detecting jailbreak prompts.
@@ 4 Proposed Method
In this section, we introduce our Eeg-Defender in detail. The overview of our framework is illustrated as Figure 3. Based on our observation that shallow layers can distinguish jailbreak prompts, we build classifiers through the transformers.
4.1 Early Exit Generation and Classifiers
We primarily develop the Eeg-Defender framework by three key steps in the following.
@@ 4.1 Early Exit Generation and Classifiers
Step II. Training Classifiers. We collect the embeddings from each layer of the LLM for prompts by generating the first token. Assuming that the LLM has $n$ layers in total, the embedding of a prompt $p_{i}$ could be represented as $E_{i}=\{e_{i1},e_{i2},\ldots,e_{in}\}$. Given the relatively small number of rejected prompts, we choose to implement prototype classifiers in our framework. The prototype $g_{ki}$ of class $k$ is computed by the mean embedding within this class Snell et al. (2017). Let $P^{{}^{\prime}}_{k}$ denote the set of samples of class $k$ in set $P^{{}^{\prime}}$. At the $i$th layer, $g_{ki}$ is represented by:
@@
$c_{i}=\arg\min_{k}d(e_{i},g_{ki})$
where $d$ represents the cosine distance as below:
$d(e_{i},g_{ki})=1-\frac{e_{i}\cdot g_{ki}}{\|e_{i}\|\|g_{ki}\|}$
Step III. Safe Generation. We use the classifiers trained in Step II to classify prompts. Based on our observations, classifiers in the early layers demonstrate higher accuracy in detecting jailbreak prompts. Consequently, the Eeg framework maintains a cumulative positive counter, referred to as the Harmfulness score, which tracks the total occurrences of positive classifications (i.e., prompts identified as harmful) by the classifier. Two hyper-parameters, $\alpha$ and $t$, control the shallow layer usage ratio and the harmfulness score threshold, respectively. Suppose the given LLM has a total of $n$ transformer layers and generates the response $x_{s+1:}$. The output of the LLM with Eeg-Defender, $x^{{}^{\prime}}_{s+1:}$ can be accessed by:
$x^{{}^{\prime}}_{s+1:}=\begin{cases}\text{Refuse to answer,}&\text{if }\sum_{i=1}^{\lfloor\alpha\times n\rfloor}c_{i}>t\\ x_{s+1:},&\text{otherwise}\end{cases}$
4.2 Eeg-Defender
Based on the classifier and configuration set in Section 4.1, Eeg-Defender can be integrated with any transformer-based LLM by monitoring the internal representation of the model. When a user inputs a prompt, Eeg-Defender calculates the harmfulness score using embeddings starting from the first layer to the $\lfloor\alpha\times n\rfloor$th layer before generating the first token. If the harmfulness score (i.e., cumulative positive count) reaches the threshold $t$, the LLM can immediately halt generation and output a standard refusal response. Essentially, Eeg-Defender evaluates the internal representations of prompts without requiring additional fine-tuning or retraining of the original model, making it a plug-and-play component for any LLM.

### v2

# Defending against Jailbreak through Early Exit Generation of Large Language Models

@@ Abstract
Large Language Models (LLMs) are increasingly attracting attention in various applications. Nonetheless, there is a growing concern as some users attempt to exploit these models for malicious purposes, including the synthesis of controlled substances and the propagation of disinformation. In an effort to mitigate such risks, the concept of "Alignment" technology has been developed. However, recent studies indicate that this alignment can be undermined using sophisticated prompt engineering or adversarial suffixes, a technique known as "Jailbreak." Our research takes cues from the human-like generate process of LLMs. We identify that while jailbreaking prompts may yield output logits similar to benign prompts, their initial embeddings within the model’s latent space tend to be more analogous to those of malicious prompts. Leveraging this finding, we propose utilizing the early transformer outputs of LLMs as a means to detect malicious inputs, and terminate the generation immediately. We introduce a simple yet significant defense approach called Eeg-Defender for LLMs. We conduct comprehensive experiments on ten jailbreak methods across three models. Our results demonstrate that Eeg-Defender is capable of reducing the Attack Success Rate (ASR) by a significant margin, roughly 85% in comparison with 50% for the present SOTAs, with minimal impact on the utility of LLMs.
Warning: this paper may contain offensive prompts and outputs.
Keywords:
Jailbreak Detection LLM Alignment
@@ 1 Introduction
Large Language Models (LLMs) are garnering unprecedented attention and application in the field of artificial intelligence, with chatbots such as ChatGPT [1] and Llama [32] standing out as notable examples. However, an inherent challenge arises because these models could generate inappropriate and potentially harmful content, including biased, unlawful, pornographic, and fraudulent material [38]. To mitigate such risks and to steer LLM-generated responses away from these issues, researchers have innovated a series of alignment algorithms [24, 36, 30]. Through the implementation of these algorithms, chatbots have been empowered to discern and tactfully refuse to generate outputs in response to prompts that naively seek to elicit potentially harmful content.
(a) Model Embedding visualization.
@@
More recently, it has been however discovered that well-designed jailbreak prompts can circumvent such alignment, posing new challenges for building stricter safety barriers [48, 18, 35]. Meanwhile, efforts to defend against jailbreaks are ongoing. Prompt-based methods [44, 40, 12, 37, 10] approach defense by manipulating or detecting user prompts. However, these methods are impractical since they degrade significantly in utility [41]. As a result, researchers turn to decoding-based defense methods [27, 3, 41, 46]. Instead of directly accessing prompts, decoding-based defense methods leverage the model’s internal properties. Since these methods can maintain high model functionality, decoding-based defense methods have shown promise in defending against jailbreak attacks.
@@
Unfortunately, current decoding-based defense technologies are insufficient. Studies show that present defense methods could only reduce the Attack Success Rate (ASR) by around 50% against jailbreak prompts [42]. Approaches like RA-LLM [3] and Smooth-LLM [27] propose generating responses multiple times with random dropouts to defend against character-sensitive adversarial suffix attacks. However, they are less effective against prompt crafting attacks, which typically involve character-insensitive prompts. SafeDecoding [41] aims to increase the likelihood of disclaimer generation artificially, but in practice, it fails to effectively reduce ASR in models with stronger safety barriers.
@@
In response to the drawbacks of existing decoding-based defense methods, we revisit the functions of different layers in LLMs. The method mentioned in [31] reveals that the initial layers specialize in triggering specific tasks. The middle layers act as repositories of knowledge and shape the emotional tone of the output [47, 45]. Subsequent layers are where the refinement of the language output occurs [7]. Given that language only affects how we deliver, but not the semantics of expression [9], we postulate that LLMs process all prompts similarly when recognizing functions in the initial layers and accessing stored knowledge in the middle layers.
@@
To validate our postulation, we conduct a series of analyses. First, our results in Section 3.2 demonstrate that the classifiers trained on the initial layers achieve over 80% accuracy in detecting fail-to-refuse harmful prompts. More intuitively, as illustrated in Figure 1(a), our empirical visualization shows that starting from the early layers of models (e.g., layer 6 and layer 8), embeddings of jailbreak prompts aligned with harmful prompts. In the middle layers (e.g., layer 12), where LLMs retrieve information, jailbreak embeddings shift towards benign embeddings slightly, and by the later layers (e.g., layers 28 and 32), they become increasingly aligned with benign embeddings. Ultimately, the jailbreak embeddings are either distributed throughout the space (as seen with Llama2) or distributed with the decision boundary (as seen with Vicuna and Guanaco), complicating the model in recognizing jailbreak status.
@@
Remarkably, the process by which large language models generate responses closely mirrors how humans organize language. To structure language output, humans first form an idea [25], then draw upon experiences and memories [5, 34]. Finally, language serves as a conduit for conveying information [2, 9]. As such, we argue that the focus may be placed on the early or intermediate layers rather than the latter or even final layers, which are overemphasized by current methods.
@@
  • Defend jailbreak through early exit. We propose Eeg-Defender. Eeg-Defender reduces Attack Success Rate (ASR) by approximately 85% against existing jailbreak methods, with near-zero computational cost.
@@ 2.2 LLM Jailbreak
Prompt Crafting. The study in [35] found that LLMs are often vulnerable to jailbreaks due to competing objectives and mismatched generalizations. They proposed 30 jailbreak methods to elicit harmful responses from GPT and Claude. To alleviate the burden of manually designing jailbreak prompts, recent studies [43, 19, 4] have introduced automated frameworks for generating effective attacks against LLMs.
Token Optimizing. In a white-box setting, attackers have access to the gradients of LLMs, allowing them to optimize prompts to increase the likelihood of generating affirmative responses. GCG attack [48] achieved jailbreak by optimizing an adversarial suffix to minimize the loss of the desired prefix of outputting. The AutoDAN attack [18]. constructed prompts that can pass perplexity testing Additionally, ICL attack [26] combined In-Context Learning with model gradients to distract the model’s attention and generate harmful content.
@@ 2.3 Jailbreak Defense
Prompt-based Defense. Detecting harmful content at the prompt level is an effective preventive strategy. To this end, Llama Guard [10], OpenAI [23], and Perspective [13] provide APIs for content moderation. Beyond detection, prompt manipulation has also been explored: Zhang et al. [44] appended prompts with safety-oriented instructions, while Self-Reminders [40] leveraged psychological cues via system prompts to encourage responsible responses. Jain et al. [12] proposed three strategies—perplexity detection, paraphrasing, and reorganization—but noted high false positive rates, limiting practical deployment.
Decoding-based Defense. Some jailbreak prompts can be highly sensitive to character-level changes. Therefore, introducing random perturbations and dropouts can help mitigate attack effects [27]. RA-LLM [3] leverages the inherent robustness of LLMs and applies Monte Carlo sampling with dropout as a defense strategy. SafeDecoding [41] revealed that safety disclaimers often remain among the top tokens in the outputs generated by jailbreak prompts. They proposed amplifying these safety token probabilities to reduce the risk of jailbreaks. Besides, Layer Editing [46] identified several safety-critical layers within LLMs and re-aligned these layers to improve overall safety. Overall, these defense methods effectively balance utility and safety, but their effectiveness diminishes with models that have stronger safety barriers.
@@ 2.4 Language Production
One of the most widely accepted theories about how language is organized in humans is Piaget’s theory, which suggests that thought forms first, and then language develops [25]. When individuals have a concept in mind, they draw upon their memories [5] and personal experiences [34, 28]. Conversely, language is optimized for communication, where people use signs to express and share their thoughts with others; this system of signs has gradually evolved into complex languages [2]. In summary, language is often seen as a bridge between communication and cognition in humans, with ideas forming first and language being structured based on memories and experiences.
Our work is inspired by the process of language production, a phenomenon also reflected in LLMs. After receiving a prompt, the LLM first identifies the purpose of the prompt and triggers a function within the model [31]. Then, it accesses and processes stored information [21] and manages emotional tone [45, 47] for prompts in the early and middle layers. Several studies found that by truncating [7], skipping [6], and pruning [20] some deeper layers, models can respond faster while maintaining correctness. This observation reveals that later layers are responsible for organizing languages. Due to the shared semantic similarities between jailbreak and harmful prompts, we believe that LLMs tend to perform similarly when identifying functions and accessing information.
@@ 3 A Closer Look into Jailbreak
Although concurrent work demonstrates that well-aligned LLMs can effectively distinguish between benign and harmful prompts within the model’s latent space [16], the mechanisms behind jailbreaks remain under debate. To gain a deeper understanding of jailbreak, we further investigate the representation of prompts.
Motivated by the human-like generation process of the language model and the observation that well-aligned LLMs can reject malicious and some jailbreak prompts, our aim is to understand how jailbreak prompts manage to bypass safety barriers. Previous attack methods [48, 35] suggest that the first token of response influenced the overall responses. Rejection responses always start with an apology or a disclaimer, while helpful responses to benign prompts typically begin with an affirmation. Given that jailbreak prompts share semantic similarities with harmful prompts but resemble benign prompts in their response patterns, we first conjecture that jailbreak embeddings progressively transit from harmful to benign as the layers go deeper.
@@ 3.1 Embedding of Jailbreak: A Toy Example
We conduct a toy example to examine how jailbreak prompts are positioned in the embedding space. We collected 60 benign prompts from Alpaca Eval [15], and 60 harmful prompts from AdvBench [48]. Then, we evaluated 60 prompts generated by GCG [48], AutoDAN [18], GPTFuzz [43], and Tap [19], all of which are effective at jailbreak models. As Figure 1(a) shows, in the final layer, the harmful prompts and benign prompts embedding are linearly separable after PCA, with jailbreak embeddings positioned between them, making detection and defense against jailbreaks more challenging. However, we found that in the earlier layers of LLMs (e.g., layer 6), embeddings for benign and harmful prompts are clearly separated, with jailbreak embeddings more closely aligned with harmful prompts. As we move to deeper layers, although benign and harmful embeddings remain distinct, jailbreak embeddings incline toward the center of benign embeddings. With this intriguing phenomenon, we also hypothesize that early and middle layers of transformers inherently possess the ability to discern jailbreak prompts.
@@ 3.2 Shallow Layers can Distinguish Jailbreak
To simulate real-world chatbot applications, we adapted the toxic-chat training dataset [17] to validate our hypothesis. The dataset includes 5,082 user prompts from the Vicuna online demo, with 384 identified as harmful. Specifically, we re-evaluated the harmful prompts in the dataset. We identified 302 harmful prompts for Llama2 and 140 harmful prompts for Vicuna that can successfully reject. Second, we collected the embedding of all layers for benign prompts and rejected harmful prompts then trained 32 MLP classifiers as well as 32 prototype classifiers corresponding to the output of each layer, respectively. We use these two sets of classifiers to identify jailbreak prompts that the model cannot reject.
Vicuna
Llama
Figure 2: Accuracy by layers.
As shown in Figure 2, classifiers trained from the early layers perform much better than those from the later layers. The accuracy in distinguishing jailbreak prompts exceeds 80% for both models up to the twelfth layer, strongly supporting our second hypothesis. This indicates that we should likely focus on the early and intermediate layer space rather than the output space.
@@ 4 Proposed Method
In this section, we introduce our Eeg-Defender in detail. The overview of our framework is illustrated as Figure 3. We primarily develop the Eeg-Defender framework by three key steps in the following.
@@
Step II. Training Classifiers. We collect the embeddings from each layer of the LLM for prompts by generating the first token. Assuming that the LLM has $n$ layers in total, the embedding of a prompt $p_{i}$ could be represented as $E_{i}=\{e_{i1},e_{i2},\ldots,e_{in}\}$. Given the relatively small number of rejected prompts, we choose prototype classifiers in our framework. The prototype $g_{ki}$ of class $k$ is computed by the mean embedding within this class [29]. Let $P^{{}^{\prime}}_{k}$ denote the set of samples of class $k$ in set $P^{{}^{\prime}}$. At the $i$th layer, $g_{ki}$ is represented by:
@@
$c_{i}=\arg\min_{k}d(e_{i},g_{ki})\;\text{,where}\;d(e_{i},g_{ki})=1-\frac{e_{i}\cdot g_{ki}}{\|e_{i}\|\|g_{ki}\|}$
Step III. Safe Generation. We use the classifiers trained in Step II to classify prompts. Based on our observations, classifiers in the early layers demonstrate higher accuracy in detecting jailbreak prompts. Consequently, the Eeg-Defender maintains a cumulative positive counter, referred to as the Harmfulness score, which tracks the total occurrences of positive classifications (i.e., prompts identified as harmful) by the classifier. Two hyper-parameters, $\alpha$ and $t$, control the shallow layer usage ratio and the harmfulness score threshold, respectively.